# A Method for the Architecture of a Medical Vertical Large Language Model Based on Deepseek R1 Mingda Zhang1, Jianglong Qin1,2*

### v1

@@ Abstract
In recent years, despite foundation models like DeepSeek-R1 and ChatGPT demonstrating significant capabilities in general tasks, professional knowledge barriers, computational resource requirements, and deployment environment limitations have severely hindered their application in actual medical scenarios. Addressing these challenges, this paper proposes an efficient lightweight medical vertical large language model architecture method, systematically solving the lightweight problem of medical large models from three dimensions: knowledge acquisition, model compression, and computational optimization. At the knowledge acquisition level, a knowledge transfer pipeline is designed from the fine-tuned DeepSeek-R1-Distill-70B teacher model to the DeepSeek-R1-Distill-7B student model, and Low-Rank Adaptation (LoRA) technology is adopted to precisely adjust key attention layers. At the model compression level, compression techniques including 4-bit weight quantization are implemented while preserving the core representation ability for medical reasoning. At the computational optimization level, inference optimization techniques such as Flash Attention acceleration and continuous batching are integrated, and a professional prompt template system is constructed to adapt to different types of medical problems. Experimental results on medical question-answering datasets show that the method proposed in this paper maintains professional accuracy while reducing memory consumption by 64.7% and inference latency by 12.4%, providing an effective solution for the application of medical large models in resource-constrained environments such as edge computing devices.

### v2

@@ Abstract
Despite significant advances in foundation models like DeepSeek-R1 and ChatGPT, their deployment in medical settings faces critical challenges including computational requirements and professional knowledge barriers. This paper presents an efficient lightweight medical large language model architecture that systematically addresses these challenges through three-dimensional optimization: knowledge acquisition, model compression, and computational enhancement. We design a knowledge transfer pipeline from DeepSeek-R1-Distill-70B to DeepSeek-R1-Distill-7B using Low-Rank Adaptation (LoRA) for precise medical knowledge retention. Through 4-bit quantization and mixed-precision strategies, we achieve substantial model compression while preserving medical reasoning capabilities. The inference framework incorporates Flash Attention acceleration and continuous batching, complemented by specialized prompt templates for diverse medical queries. Experimental evaluation on medical benchmarks demonstrates that our approach maintains 92.1% accuracy on USMLE examinations while reducing memory consumption by 64.7% and inference latency by 12.4% compared to baseline models. This work provides a practical solution for deploying advanced language models in resource-constrained medical environments, enabling broader accessibility of AI-assisted healthcare.